# SoK: A Classification for AI-driven Personalized Privacy Assistants

### v2

@@ Abstract.
To help users make privacy-related decisions, personalized privacy assistants based on AI technology have been developed in recent years. These AI-driven Personalized Privacy Assistants (AI-driven PPAs) can reap significant benefits for users, who may otherwise struggle to make decisions regarding their personal data in environments saturated with privacy-related decision requests. However, no study systematically inquired about the features of these AI-driven PPAs, their underlying technologies, or the accuracy of their decisions. To fill this gap, we present a Systematization of Knowledge (SoK) to map the existing solutions found in the scientific literature. We screened 1697 unique research papers over the last decade (2013-2023), constructing a classification from 39 included papers. As a result, this SoK reviews several aspects of existing research on AI-driven PPAs in terms of types of publications, contributions, methodological quality, and other quantitative insights. Furthermore, we provide a comprehensive classification for AI-driven PPAs, delving into their architectural choices, system contexts, types of AI used, data sources, types of decisions, and control over decisions, among other facets. Based on our SoK, we further underline the research gaps and challenges and formulate recommendations for the design and development of AI-driven PPAs as well as avenues for future research.
@@
privacy assistant, privacy, data protection, artificial intelligence, machine learning, systematic review
journalyear:
1. Introduction
@@
As the world becomes increasingly digitalized, people are faced with a higher number of decisions related to their privacy. We surround ourselves daily with several apps and websites, and the number of smart gadgets and Internet of Things (IoT) devices continues to grow (noa, ). Furthermore, to enforce the individuals’ rights to informational self-determination and comply with privacy laws such as the General Data Protection Regulation (GDPR) (European Commission, 2016), software systems regularly require us to make privacy-related decisions regarding our personal data: Do you grant this permission? Do you want to accept the cookies? Should this sensor be left on when you host friends? Consequently, the cognitive burden increases, leaving users in disarray, tired, and unable to decide in their best interests (Choi et al., ).
During the last decade, researchers have been building privacy assistants to alleviate this burden and support users in their decisions (see the patent on Personalized Privacy Assistant registered in 2023 in the US by Sadeh et al. (2021)). With the progress made in Artificial Intelligence (AI), it is no surprise that some of these assistants leverage this technology, notably enabling more personalized support. However, the extent to which AI drives these Personalized Privacy Assistants (AI-driven PPAs), their efficiency, privacy-friendliness, functioning, and eventual addressing of legal requirements remains unclear. In fact, to the best of our knowledge, there have been no surveys or systematic reviews on the topic of AI-driven PPAs. This lack of systematization of knowledge prevents other researchers from identifying existing gaps in the field and efficiently addressing the challenges.
@@
To address this lack of coherence, provide a common vocabulary, and better compare and categorize the different AI-driven PPA solutions, we propose a Systematization of Knowledge (SoK) of the last decade of research. In doing so, we aim to draw insights and lessons for future assistants and to formulate better recommendations for research, design, and development of AI-driven PPAs. Formulated otherwise, we tackle the following Research Questions (RQs):
@@
Here, we understand agents and assistants in a broad sense (any logical entity able to support users, including unimplemented theoretical models, see our selection criteria in Table 1); AI in a generic sense as well (see Section 2.3); and privacy decisions as individual decisions regarding one’s personal information management (see Section 2.2).
@@
To address our RQs, we performed a Systematic Literature Review (SLR) on research papers that provided technical solutions, published between 2013 and 2023 in peer-reviewed venues, and a further snowballing process until early May 2024. We screened 1697 unique papers from IEEE, ACM, Scopus, and Web of Science, resulting in 39 selected papers after several rounds of snowballing. We extensively read and analyzed all the included papers, and the information extracted forms the basis of our work.
As a result of our SLR, we make the following contributions:
A Classification for AI-driven PPAs:
@@
- We propose the first classification for AI-driven PPAs, providing a common vocabulary for designers of such systems.
Data Charting & Quantification:
@@
- We charted and quantified several aspects of AI-driven PPAs based on the aforementioned classification.
Research Gaps & Challenges:
@@
- We underline the current gaps in the state of the art and highlight challenges for designing AI-driven PPAs based on our data.
Recommendations & Research Avenues:
@@
- We formulate recommendations for improving AI-driven PPAs, and propose several avenues for future research.
2. Background and Related Work
@@
As a background, this section first introduces legal concepts around privacy, then provides an overview of different types of privacy decisions for which individuals could receive support from AI-driven PPAs. Lastly, it presents AI and Machine Learning technologies that can provide the technical foundations for AI-driven PPAs.
2.1. Legal Background
2.1.1. Roles and Obligations According to the GDPR and AI Act
AI-driven PPAs are using AI techniques for processing data, including personal data, to assist users with making personalized privacy-related decisions. For the discussion of any legal requirements regarding the use of personalized privacy assistants, the question of who the data controller is for any personal data processing by the assistants will be of relevance.
If AI-driven PPAs are installed and run by users on their own devices or other servers under their control, the users will likely act as data controllers or joint controllers with other service providers. The so-called household exemption (Art. 2(2)(c) GDPR) can take effect, meaning that the GDPR (European Commission, 2016) will not apply if the user is using the assistant for private purposes on their private devices or servers under their control. If the AI-driven PPA is run not only for purely private purposes on the user’s devices or controlled servers, the data controller may be another entity different from the user (e.g., the user’s employer). Legal obligations need to be fulfilled by the controllers regarding data protection by design and default (Art. 25) of the assistants, security of data processed (Art. 32), implementing data subject rights, including the data subject’s rights to transparency (Art. 13–15), their rights to object to profiling (Art. 21), and the right not to be subject to a decision based solely on automated processing, including profiling (Art. 22).
In addition, legal obligations according to the EU AI Act (European Commission, 2024) may also have to be considered for the producer and also by the deployers of AI-driven privacy assistants, including requirements for risk management (Art. 9), transparency (Art. 13, 50), robustness, security, accuracy (Art. 15). These obligations however mostly apply if AI-driven PPAs could be classified as “high-risk” AI systems. This should, however, seldom be the case, especially as AI-driven PPAs are typically used for users’ own privacy management, which should typically not interfere with the fundamental rights of others. Exceptions could, however, be AI-driven PPAs that are, for example, used for setting permissions for safety-critical applications impacting the safety of the users or others.
2.1.2. Legal Requirements for Transparency.
In cases where the data controllers of the AI-driven PPAs are not the data subjects themselves, the controllers should provide the data subjects with privacy policy information ex-ante at the time when data are obtained from them according to Art. 13 GDPR, and ex-post through the right to access granted in Art. 15 GDPR. This also should include information about purposes of processing, data categories concerned, but also information about the logic involved and significance, and envisioned consequences of automated decision-making and profiling performed by the AI-driven PPAs.
The AI Act also includes obligations for transparency for the producers and deployers of limited-risk and high-risk AI systems (Art. 50). While the providers of limited-risk AI systems have to mainly ensure that humans are informed that AI systems are used, high-risk AI systems require that further clear, comprehensible and adequate information is given to the deployer (Art. 13), traceability of results via logging (Art. 12) and appropriate human oversight (Art. 14).
2.1.3. Legal Requirements for Consent.
Art. 4 (11) of the GDPR defines ‘Consent’ of data subjects as any freely given, specific, informed, and unambiguous indication of the data subject’s wishes by which they, by a statement or by a clear affirmative action, signifies agreement to the processing of personal data relating to them. A valid consent has thus to fulfill several conditions. Namely, it needs to be:
  • Freely given, i.e., the data subject needs to have free choices – this is usually not the case if there is an imbalance of power in the relation between the data subject and the data controller. Furthermore, there should be no negative consequences if consent is not given. Moreover, consent may not be bundled as a non-negotiable part of terms and conditions.
  • Specific, which means that consent must be given for one or more specific purposes and that the data subject must have a choice in relation to them – i.e., separate opt-in is needed for each purpose.
  • Informed, which means that the data subject has to be informed about certain elements that are crucial to making a choice. This includes information about the controller’s identity, the data processing purposes, the type of data, the right to withdraw consent, any use for decisions based solely on automated processing, and risks of data transfers to third countries.
  • Moreover, a confirming statement or affirmative action is needed for a valid consent and requires that the data subject has taken a deliberate action to consent. Therefore, silence, pre-ticked boxes, or inactivity should not constitute consent (Recital 32 GDPR). It also means that consent cannot be fully automated.
2.2. Privacy Decisions
@@
Among the most notable definitions, Westin (Westin, ) has defined privacy as the right to informational self-determination, meaning that individuals should have the right to decide for themselves when, how, and what information about them is communicated to others. As mentioned, in the EU, the GDPR emphasizes that individuals should have control of their personal data (Recital 7), and thus should be empowered to make decisions about their data as one prerequisite for exercising such control. Delving deeper into this notion of privacy decisions, we further elaborate on this concept in the following subsections.
2.2.1. Individual Privacy Decisions Regulated by Privacy Laws
@@
Some privacy decisions individuals can make to exercise control over their data are regulated under the GDPR and other privacy laws. These decisions notably include, but are not limited to, the decisions to grant or to withdraw consent to data collection and processing.
@@
Indeed, the GDPR and most other privacy laws regulate decisions to exercise data subject rights granted by the respective laws. For instance, according to Art. 15-22 GDPR, data subjects have the rights to access data, request rectification or deletion of data, export data, and object to direct marketing and profiling. Data subjects can also object in cases where the legal ground for the processing is public interest or legitimate interest, or exercise their right not to be subject to automated decision-making.
2.2.2. Further Types of Privacy Decisions
@@
Moreover, privacy decisions encompass privacy permission (or access control rights) settings, which grant others certain rights for using their data and are, for instance, typically used for permission systems of mobile phone operating systems, such as Android or iOS. Setting privacy permissions on mobile operating systems often requires consent at installation or during runtime. However, instead of consent, other legal grounds – such as a contract (Art. 6 (1)(b) GDPR) –, can be used, e.g., for a banking app to forward account information when transferring money (Art. 29 Working Party, 2013). Let us also note the peculiar case of Global Privacy Control (GPC), a unary signal that permits or prohibits third-party tracking on the browser (Human et al., ). Due to its enforceability under the California Consumer Privacy Act (CCPA), it is regulated by a privacy law but is technically more akin to a privacy permission.
@@
Additionally, some privacy-enhancing technologies and protocols allow users to decide and set privacy preferences, which are simply indications of the users’ privacy wishes of how their data should be used without actually granting any rights to others, and thus without legal mandate. Privacy preferences have, for instance, been used earlier by the Platform for Privacy Preferences (P3P) (Cranor et al., 2002) or Do Not Track (DNT) as an example for signals that can be set manually in browser settings for allowing users to specify their privacy choices.
2.3. AI for Decision-making
@@
AI is a generic term for various strategies and techniques enabling computers and machines to simulate human intelligence and problem-solving capabilities (Russell and Norvig, ). Machine learning (ML) is a field of AI (we subsume the former under the latter in the rest of the document) that develops and studies statistical algorithms and models, draws inferences from patterns in data, and learns and adapts without following explicit instructions. AI-powered tools can particularly lighten the user’s cognitive load and thereby improve their decision-making, e.g., by decision support, augmentation, or automation.
@@
While there are different ways to categorize AI systems, we refer in the present work to the survey paper on eXplainable AI (XAI) by Arrieta et al. . They distinguish between transparent models and those requiring post-hoc explainability (non-inherently transparent). We leverage this reference because AI-supported decision-making must be explained under specific circumstances according to the GDPR and the AI Act (Panigutti et al., ).
@@
In their words: “A model is considered to be transparent if by itself it is understandable.” (Arrieta et al., ) Such models include linear regression, decision trees, k-nearest neighbors, rule-based learning, general additive models, and Bayesian models. Nonetheless, models that are not deemed intrinsically transparent can be made explainable through the use of post-hoc techniques. Neural networks (especially deep and convoluted) and Support Vector Machines (SVM) typically fall under this category, as well as reinforcement learning (Puiutta and Veith, ).
3. Methodology
@@
This SoK study adopts the widely known methodology for systematic literature reviews (SLRs) proposed by Kitchenham (2004). The SLR methodology offers us a well-defined and rigorous sequence of methodological steps consisting of three main phases: (1) planning, (2) conducting, and (3) reporting the review. A SLR Protocol that describes the entire research process has been written for this study (a summary version of which can be found in Appendix A.1). Furthermore, we make our research data openly available in an anonymised GitHub repository  [^1] https://anonymous.4open.science/r/SoK_AI_PPA-E29F for reproducibility. Our material comprises the citation files of each query, the Data Extraction Forms (DEFs) of the selected papers, and the charting spreadsheet used to compile all our data. Thus, due to page constraints, we refer readers to these documents for methodological details.
3.1. Planning the Review
@@
During the planning phase, our first activity was determining the need for this SLR. Several databases were searched to verify if any surveys or reviews had been conducted on AI-driven PPAs. Search terms such as privacy, data protection, assistant, agent, artificial intelligence, and machine learning were used. However, we could not identify any survey or systematic reviews on the topic, reassuring the need for an SLR.
The research questions, presented in Section 1, guided the remaining phases of this SLR with respect to the search process, selection criteria, and data synthesis.
Figure 1. Sankey chart of the selection process.
3.2. Conducting the Review
3.2.1. Search Strategy

### v4

@@ Abstract
In recent years, several personalized assistants based on AI have been researched and developed to help users make privacy-related decisions. These AI-driven Personalized Privacy Assistants (AI-driven PPAs) can provide significant benefits for users, who might otherwise struggle with making decisions about their personal data in online environments that often overload them with different privacy decision requests. So far, no studies have systematically investigated the emerging topic of AI-driven PPAs, classifying their underlying technologies, architecture and features, including decision types or the accuracy of their decisions. To fill this gap, we present a Systematic Literature Review (SLR) to map the existing solutions found in the scientific literature, which allows reasoning about existing approaches and open challenges for this research field. We screened several hundred unique research papers over the recent years (2013-2025), constructing a classification from 41 included papers. As a result, this SLR reviews several aspects of existing research on AI-driven PPAs in terms of types of publications, contributions, methodological quality, and other quantitative insights. Furthermore, we provide a comprehensive classification for AI-driven PPAs, delving into their architectural choices, system contexts, types of AI used, data sources, types of decisions, and control over decisions, among other facets. Based on our SLR, we further underline the research gaps and challenges and formulate recommendations for the design and development of AI-driven PPAs as well as avenues for future research.
Index Terms:
@@
artificial intelligence, data protection, machine learning, privacy, privacy assistant, systematic review
[…]
=-21pt
I Introduction
@@
As the world becomes increasingly digitalized, people are faced with a growing number of requests for decisions related to their online privacy. Nowadays, individuals are using several apps every day, visiting different websites, and the number of smart gadgets and Internet of Things (IoT) devices they use continues to grow [1]. Furthermore, to comply with privacy laws such as the General Data Protection Regulation (GDPR) [2], software systems frequently demand from us to make privacy-related decisions regarding our personal data: Do you grant this permission? Do you want to accept the cookies? Should this sensor be left on when you host friends? Consequently, the cognitive burden increases, leaving users in disarray, tired, and unable to decide in their best interests [3].
During the last decade, researchers have been building privacy assistants to alleviate this burden and support users in their decisions. One of the first research work in that field has resulted in the patent on a Personalized Privacy Assistants (PPAs), registered in 2023 in the United States by Sadeh et al. [4]). With the progress made in Artificial Intelligence (AI), it comes at no surprise that many of the privacy assistants proposed or developed in recent years leverage AI technology, notably to enable better personalized support.
This personalization can enhance the quality of decision support, adapted to the individuals’ needs, preferences and current context. However, the extent to which AI drives these AI-driven PPAs, their efficiency, privacy-friendliness, functionality, and how legal requirements are eventually addressed remains unclear. In fact, to the best of our knowledge, there have been no surveys or systematic reviews on the topic of AI-driven PPAs, despite the substantial amount of work published on the topic in the recent years. This lack of systematization of knowledge makes it more difficult for other researchers and developers to reason about the opportunities that current AI-driven PPAs may offer. It also makes it more difficult to identify limitations and existing gaps to be addressed, as well as open research challenges that remain for the future.
@@
For this reason, we present a Systematic Literature Review (SLR) of the body of knowledge to provide a common vocabulary and better compare, categorize and analyze the different AI-driven PPA solutions. In doing so, we aim to draw insights and lessons for future assistants and to formulate better recommendations for research, design, and development of AI-driven PPAs. Therefore, this SLR addresses the following Research Questions (RQs):
@@
Here, we consider agents and assistants in a broad sense (any logical entity able to support users, including unimplemented theoretical models, see our selection criteria in Table I); AI in a generic sense as well (see Section II-C); and, privacy decisions as the individual’s decisions regarding their personal information management (see Section II-A).
@@
To address our RQs, we performed an SLR on research papers providing technical solutions, published between 2013 and 2025 in peer-reviewed venues, including a snowballing process until early February 2025. We screened several hundred papers from IEEE, ACM, Scopus, and Web of Science, resulting in 41 selected papers after several rounds of snowballing. We extensively read and analyzed all included papers, and the information extracted forms the basis of our work.
Our SLR results in the following contributions:
@@
  • A Classification for AI-driven PPAs – We propose the first classification for AI-driven PPAs, providing a common vocabulary for designers of such systems.
@@
  • Data Charting & Quantification – We charted and quantified several aspects of AI-driven PPAs based on the aforementioned classification.
@@
  • Research Gaps & Challenges – We underline the current gaps in the state of the art and highlight challenges for designing AI-driven PPAs based on our data.
@@
  • Recommendations & Research Avenues – We formulate recommendations for improving AI-driven PPAs, and propose several avenues for future research.
In the following sections, we present the background and related work in Section II. The study’s methodology is detailed in Section III. The results and classification are organized and presented in Sections IV and V. Based on the findings, we present our discussion of research gaps and future work in Section VI. Lastly, Section VIII concludes our work.
II Background
@@
As a background, this section first provides an overview of different types of privacy decisions for which individuals could receive support from AI-driven PPAs. Then, it summarizes legal requirements for transparency that AI-driven PPAs should meet, and refers to a classification scheme of explainable AI that we are using for our classification.
When discussing legal requirements, we will primarily refer in the section and for the rest of this paper to the European Legal Framework, including the GDPR and the AI Act [5], since the study was conducted in Europe with the support of a European funding foundation. Moreover, the GDPR has been regarded as the “gold standard” for data protection with a territorial scope that goes beyond Europe, and is therefore also used as a point of reference.
II-A Privacy Decisions
@@
Among the most notable definitions, Westin [6] has defined privacy as the right to informational self-determination, meaning that individuals should have the right to decide for themselves when, how, and what information about them is communicated to others. As mentioned, in the EU, the GDPR emphasizes that individuals should have control of their personal data (Recital 7), and thus should be empowered to make decisions about their data as one prerequisite for exercising such control. Delving deeper into this notion of privacy decisions, we further elaborate on this concept in the following subsections.
II-A1 Individual Privacy Decisions Regulated by Privacy Laws
@@
Some privacy decisions individuals can make to exercise control over their data is regulated under the GDPR and other privacy laws. These decisions notably include, but are not limited to, the decisions to grant or to withdraw consent to data collection and processing. Art. 4 (11) of the GDPR defines ‘consent’ of data subjects as any freely given, specific, informed, and unambiguous indication of the data subject’s wishes by which they, by a statement or by a clear affirmative action, signifies agreement to the processing of personal data relating to them.
@@
Moreover, the GDPR and most other privacy laws regulate further decisions to exercise data subject rights granted by the respective laws. For instance, according to Art. 15-22 GDPR, data subjects have the rights to access data, request rectification or deletion of data, export data, and object to direct marketing and profiling. Data subjects can also object in cases where the legal ground for the processing is public interest or legitimate interest, or exercise their right not to be subject to automated decision-making.
II-A2 Further Types of Privacy Decisions
@@
Moreover, privacy decisions encompass privacy permission settings (or access control rights), which grant others certain rights for using their data and are, for instance, typically used for permission systems of mobile phone operating systems, such as Android or iOS. Setting privacy permissions on mobile operating systems often requires consent at installation or during runtime. However, instead of consent, other legal grounds – such as a contract (Art. 6 (1)(b) GDPR) –, can be used, e.g., for a banking app to forward account information when transferring money [7]. Let us also note the peculiar case of Global Privacy Control (GPC), a unary signal that permits or prohibits third-party tracking on the browser [8]. Due to its enforceability under the California Consumer Privacy Act (CCPA) [9], it is regulated by a privacy law but is technically more akin to a privacy permission.
@@
Additionally, some privacy-enhancing technologies and protocols allow users to decide and set privacy preferences, which are simply indications of the users’ privacy wishes of how their data should be used without actually granting any rights to others, and thus without legal mandate. Privacy preferences have, for instance, been used earlier by the Platform for Privacy Preferences (P3P) [10] or Do Not Track (DNT) [11], as an example for signals that can be set manually in browser settings for allowing users to specify their privacy choices.
II-B Requirements for Transparency.
Transparency of data processing is an important prerequisite for users for making well-informed decisions, and should therefore be provided by any privacy assistants that should support users in decision-making. In cases where the data controllers of the AI-driven PPAs are not the data subjects themselves, the controllers should provide the data subjects with privacy policy information ex-ante at the time when data is obtained from them according to Art. 13 GDPR, and ex-post through the right to access granted in Art. 15 GDPR. This should particularly include information about purposes of processing, data categories concerned, but also information about the logic involved and significance, and envisioned consequences of automated decision-making and profiling performed by the AI-driven PPAs.
The EU AI Act also includes obligations for transparency for the producers and deployers of limited-risk and high-risk AI systems (Art. 50). While the providers of limited-risk AI systems have to mainly ensure that humans are informed that AI systems are used, high-risk AI systems require that further clear, comprehensible and adequate information is given to the deployer (Art. 13), traceability of results via logging (Art. 12) and appropriate human oversight (Art. 14). However, AI-driven PPAs are typically not in the high-risk category, since they are used for users’ own personal privacy management, which should typically not interfere with the fundamental rights of others. Exceptions could, however, be AI-driven PPAs that are, for example, used for setting permissions for safety-critical applications impacting the safety of the users or others.
Also, Ethics Guidelines for Trustworthy AI, promoted by the EU Commission, [^1] https://digital-strategy.ec.europa.eu/en/library/ethics-guidelines-trustworthy-ai emphasize the requirement for transparency and explainability for AI systems to be deemed trustworthy.
II-C AI for Decision-making
@@
AI is a generic term for various strategies and techniques enabling computers and machines to simulate human intelligence and problem-solving capabilities [12]. Machine learning (ML) is a field of AI (we subsume the former under the latter in the rest of the document) that develops and studies statistical algorithms and models, draws inferences from patterns in data, and learns and adapts without following explicit instructions. AI-powered tools can particularly lighten the user’s cognitive load and thereby improve their decision-making, e.g., by decision support, augmentation, or automation.
@@
While there are different ways to categorize AI systems, we refer in the present work to the survey paper on eXplainable AI (XAI) by Arrieta et al. [13]. They distinguish between transparent models and those requiring post-hoc explainability (called non-inherently transparent). We use this reference because AI-supported decisions must be explained under specific circumstances according to the GDPR and the AI Act [14].
@@
In their words: “A model is considered to be transparent if by itself it is understandable.” [13]. Such models include linear regression, decision trees, k-nearest neighbors, rule-based learning, general additive models, and Bayesian models. Nonetheless, models that are not deemed intrinsically transparent can be made explainable through the use of post-hoc techniques. Neural networks (especially deep and convoluted) and Support Vector Machines (SVM) typically fall under this category, as well as reinforcement learning [15].
III Methodology
@@
This study adopts the widely known methodology for systematic literature reviews (SLRs) proposed by Kitchenham [16]. The SLR methodology offers us a well-defined and rigorous sequence of methodological steps consisting of three main phases: (1) planning, (2) conducting, and (3) reporting the review. An SLR Protocol that describes the entire research process has been written for this study. Furthermore, we make our research data openly available in a GitHub repository for reproducibility. [^2] https://github.com/Victor-Morel/SLR_AI_PPA Our material comprises the citation files of each query, the Data Extraction Forms (DEFs) of the selected papers, and the charting spreadsheet used to compile our data. We refer to these documents for methodological details.
III-A Planning the Review
@@
The first activity of the planning phase was to determine the need for this SLR. Several databases were searched to verify if any surveys or reviews had been conducted on AI-driven PPAs. Search terms such as privacy, data protection, assistant, agent, artificial intelligence, and machine learning were used. However, we could not identify any survey or systematic reviews on the topic, reassuring the need for an SLR. The research questions, presented in Section I, guided the remaining phases of this SLR with respect to the search process, selection criteria, and data synthesis.
III-B Conducting the Review
III-B1 Search Strategy